Work, energy and power(2)vector:addition of vector resolution (3)speed and acceleration (4 )rectilinear acceleration

Grade:10

1 Answers

Arun
6 years ago
Dear student
 

work is related to the amount of energy transferred to or from a system by a force. It is a scalar-valued quantity with SI units of Joule.

Work can be represented in a number of ways. For the case where a body is moving in a steady direction, the work done by a constant force F acting parallel to the displacement {\displaystyle \Delta x} is defined as

{\displaystyle W_{F}=F~\Delta x\,\!.}

When the force is not acting parallel to the body's direction of movement, the work done is defined as a dot product of the force and the displacement,

 

Power-

Power is defined to be the rate at which work is performed, or the derivative of work over time. The SI unit for power is the watt.

{\displaystyle P={\frac {\mathrm {d} E}{\mathrm {d} t}}={\frac {\mathrm {d} W}{\mathrm {d} t}}}

Average power is the average amount of work done per unit of time. Thus instantaneous power is the limiting power of the average power as Δt approaches zero

 

Energy-

Energy is stored work. It has the same units as work, the Joule (J).

 

 

Regards

Arun (askIITians forum expert)